OpenStudy (anonymous):

Solve the equation using the Zero-Product Property. (x-2)(x+7)=0 A. 2, -7 B. 2, 7 C. –1,1 D. –2, -7

OpenStudy (anonymous):

any guesses?

OpenStudy (anonymous):

i think its B?

OpenStudy (anonymous):

set \(x-2=0\) solve for \(x\) then set \(x+7=0\) and solve for \(x\)

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

set each factor equal to zero and solve if \(x-2=0\) then \(x=2\) and if \(x+7=0\) then \(x=-7\)